# Goal-Oriented Semantic Communication for Wireless Image Transmission via Stable Diffusion

### v1

@@ III-C PSD Denoiser under Unknown Channel
Notably, Algorithm 1 is only applicable when the instantaneous channel gain $\bm{h}$ is known, and hence cannot be directly used for the scenarios (e.g., Massive MIMO communications systems) with imperfect estimation of $\bm{h}$. To solve this issue, we propose a parallel SD-GSC (PSD-GSC) to jointly estimate the channel gain and remove the noise.
@@ III-C1 Forward Process
$p(\bm{z},\bm{h}|\bm{\bm{z}^{\prime}})=p(\bm{\bm{z}^{\prime}}|\bm{z},\bm{h})p(\bm{z})p(\bm{h}).$ (19)
Therefore, we can train two separate forward processes for $\bm{z}$ and $\bm{h}$, respectively. Similar to (12) and (17), the score network of $\bm{h}$ and $E[\bm{h}_{0}|\bm{h}_{t}]=\hat{\bm{h}}_{0}(t)$ can be estimated using
@@
$\hat{\bm{h}}_{0}(t)=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\vartheta}(\bm{z}_{t},t)),$ (21)
@@ III-C2 Reverse Process
Similar to the reverse process in (13), the two reverse processes of PSD-GSC can be expressed as
@@
Correspondingly, the above two reverse processes can be discretized in a similar way as (14) and expressed as
$\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+$ $\beta_{t}[\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$ (24)
$+\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)])+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}),$
@@
$\bm{h}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}+$ $\beta_{t}[\nabla_{\bm{h}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$ (25)
$+\bm{s}_{\vartheta}\left(\bm{\bm{h}}_{t},t\right)])+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}).$
@@
The detailed process of our proposed PSD denoiser for GSC (PSD-GSC) under unknown channel is presented in Algorithm 2.
0:  Training dataset $\mathcal{X}$, learning rate $\gamma$, time steps $T$, hyperparameter $\left\{\beta_{1},\cdots,\beta_{T}\right\}$, $\zeta_{t}^{\theta}$, $\zeta_{t}^{\vartheta}$ and $\phi$
0:  Reconstructed image $\tilde{\bm{x}}$.
1:  Train semantic autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ using steps 1-8 in Algorithm 1, then freeze the parameters.
// Training semantic denoiser
2:  Train denoising estimator $\epsilon_{\theta}$ and $\epsilon_{\vartheta}$ for $\bm{z}$ and $\bm{h}$ following steps 10-15 in Algorithm 1, respectively.
// Denosing process
3:  Sample $\bm{z}_{T},\bm{h}_{T}\sim\mathcal{N}(0,I)$.
4:  for denosing step $t=T,\ldots,1$ do
5:     $\hat{\epsilon}_{\theta}\leftarrow\epsilon_{\theta}(\bm{z}_{t},t)$, $\hat{\epsilon}_{\vartheta}\leftarrow\epsilon_{\vartheta}(\bm{z}_{t},t)$.
6:     Estimate $\hat{\bm{z}}_{t}$ and $\hat{\bm{h}}_{t}$ using (17) and (21).
7:     $g_{\theta}\leftarrow\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$, $g_{\vartheta}\leftarrow\nabla_{\bm{h}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$.
8:     Compute the conditional score
@@
9:     Sample $\varphi_{\theta},\varphi_{\vartheta}\sim\mathcal{N}(0,I)$.
10:     Compute $\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+\beta_{t}\bm{s}_{\theta})+\sqrt{\beta_{t}}\varphi_{\theta}$ and $\bm{h}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}+\beta_{t}\bm{s}_{\vartheta})+\sqrt{\beta_{t}}\varphi_{\vartheta}$.
11:     ${\bm{h}_{t-1}\leftarrow\bm{h}_{t-1}-\alpha(||\bm{\bm{z}^{\prime}}-\bm{h}_{t-1}*\bm{z}_{t-1}||_{2}+\phi||\bm{h}_{t-1}||)}$.
12:  end for
Algorithm 2 PSD-GSC under unknown channel
@@ IV Performance Evaluation
We implement the image transmission task under a Rayleigh fading channel, using the FFHQ dataset [14], which consists of 70,000 high-quality human face images at a resolution of 256×256 pixels, that is created as a benchmark for generative tasks. In the implementation of the semantic autoencoder, the encoder downsamples the feature dimensions by half, starting from the 2nd block, with channel dimensions specified as $[64,128,256,512]$, and then the features are embedded into a latent representation of dimension $d=4096$. Conversely, the decoder upsamples the feature dimensions by double for the first 3 blocks, with channel dimensions set to be $[512,256,128,64]$. For the semantic denoiser, we set the time step $T=1000$ and use a linear variance scheme to determine the hyperparameters of $\left\{\beta_{1},\beta_{2},\cdots,\beta_{T}\right\}$.
@@ IV-A Known Channel
Fig. 2 presents the PSNR and the FID performance across various SNRs under known channel. As expected, both PSNR and FID performances improve as the SNR increases, owing to the fact that better wireless channel conditions facilitate higher-quality image reconstruction. Interestingly, the Latent-Diff DNSC scheme outperforms the ADJSCC scheme in both PSNR and FID at all SNRs, due to diffusion model can efficiently capture the wireless channel characteristics and remove the noise from the received noisy semantic information, resulting in a more accurate reconstruction. Importantly, our proposed SD-GSC outperforms two baseline schemes in both PSNR and FID metrics. Specifically, compared to Latent-Diff DNSC and ADJSCC, the PSNR of our proposed SD-GSC achieves approximately 5 dB and 7dB improvement, and the FID of our proposed SD-GSC achieves around 16 and 30 reduction, respectively. This is because integrating the instantaneous channel gain will enable better characterization of channel gain and mitigation of distortion due to channel noise, resulting in a more controlled and guided image generation than the uncontrolled generation process of DDPM used in the Latent-Diff DNSC scheme.
@@ IV-B Unkown Channel
To validate the effectiveness of our proposed PSD-GSC for channel estimation as well as efficient image transmission, we also compare our proposed PSD-GSC with MSD-GSC and SD-GSC. Fig. 3 plots the PSNR and the FID for various SNRs under unknown channel. As expected, our proposed SD-GSC with known $h$ outperforms PSD-GSC and MSD-GSC with unknown $h$ in terms of both PSNR and FID, due to the errors in unknown channel estimation. Specifically, our proposed PSD-GSC and MSD-GSC experience degradation in PSNR by approximately 1 dB and 3 dB, and degradation in FID by 2 and 8, respectively. This is because channel estimation errors result in inaccurate channel gain inputs of the diffusion model, which decreases its denoising and high-quality image reconstruction capabilities. Importantly, PSD-GSC outperforms MSD-GSC both in PSNR and FID. This is due to the probabilistic modeling-based diffusion model used in PSD-GSC can well estimate nonlinear channels and capture the complex properties of wireless channels, for better image reconstruction.
@@ V Conclusion
In this paper, we developed a stable diffusion (SD)-based goal-oriented semantic communication framework (SD-GSC) for efficient image transmission over Rayleigh fading channels. The main goal is to reduce the bandwidth requirement while ensuring high-quality image reconstruction at the receiver. To this end, we first designed a semantic autoencoder to extract semantic information to reduce the transmitted data size while accurately reconstructing the image from the received semantic information. To combat the negative impact of wireless channel noise on the transmitted semantic information, we designed an SD-based semantic denoiser to better characterize and mitigate the channel noise, leading to significant improvements in both PSNR and FID metrics compared to the state-of-the-art baselines under known channel. For unknown channel scenarios, we further designed a parallel SD-GSC (PSD-GSC) to jointly learn the channel gain and denoise the received semantic information, showcasing the capability of our proposed PSD-GSC in channel estimation and efficient image transmission.

### v2

@@ III-C PSD Denoiser under Unknown Channel
Notably, the SD denoiser is only applicable when the instantaneous channel gain $\bm{h}$ is known, and hence cannot be directly used for the scenarios with imperfect estimation of $\bm{h}$ (e.g., Massive MIMO communications systems). To solve this issue, we propose a parallel SD (PSD) denoiser to jointly estimate the channel gain and remove the noise.
@@ III-C1 Forward Process
$p(\bm{z},\bm{h}|\bm{\bm{z}^{\prime}})\propto p(\bm{\bm{z}^{\prime}}|\bm{z},\bm{h})p(\bm{z})p(\bm{h}).$ (17)
Therefore, we can train two separate forward processes for $\bm{z}$ and $\bm{h}$, respectively. Similar to (10) and (15), the score network of $\bm{h}$ and $E[\bm{h}_{0}|\bm{h}_{t}]=\hat{\bm{h}}_{t}$ can be estimated using
@@
$\hat{\bm{h}}_{t}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\vartheta}(\bm{z}_{t},t)),$ (19)
@@ III-C2 Reverse Process
Similar to the reverse process in (11), the two reverse processes of PSD can be expressed as
@@
Correspondingly, the above two reverse processes can be discretized in a similar way to (12) and expressed as
$\bm{z}_{t-1}=$ $\frac{1}{\sqrt{\alpha_{t}}}\big{(}\bm{z}_{t}+\beta_{t}\big{[}\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})+\bm{s}_{\theta}\left(\bm{\bm{z}}_{t},t\right)\big{]}\big{)}$
$+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}),$ (22)
@@
$\bm{h}_{t-1}=$ $\frac{1}{\sqrt{\alpha_{t}}}\big{(}\bm{h}_{t}+\beta_{t}\big{[}\nabla_{\bm{h}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})+\bm{s}_{\vartheta}\left(\bm{\bm{h}}_{t},t\right)\big{]}\big{)}$
$+\sqrt{\beta_{t}}\mathcal{N}(0,\mathbf{I}).$ (23)
@@
III-D Training and Implementation
The proposed SD-based GSC architecture can be effectively trained via an end-to-end approach; however, this method may suffer from slow convergence. To mitigate this issue, we employ a step-by-step training strategy that systematically trains each component of the SD-based GSC architecture.
0:  Training dataset $\bf{X}$, time steps $T$, hyperparameters $\left\{\beta_{1},\cdots,\beta_{T}\right\}$, $\zeta_{t}^{\theta}$, $\zeta_{t}^{\vartheta}$, and $\phi$
0:  Semantic denoisers $\zeta_{t}^{\theta}$, $\zeta_{t}^{\vartheta}$
1:  Train the semantic encoder and decoder using (25).
// Training SD denoiser
2:  Freeze parameters of semantic extraction module $\mathcal{E}(\cdot)$.
3:  repeat
4:     Sample a latent representation $\bm{z}_{0}\sim p(\bm{z})$.
5:     $t\sim\textrm{Uniform}(\{1,2,\ldots,T\})$, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
6:     Take gradient descent step on $\nabla_{\theta}\left\|\epsilon_{\theta}(\bm{z}_{t},t)-\epsilon\right\|_{2}^{2}$.
7:  until Converged
// Denoising of SD denoiser
8:  Sample $\bm{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$
9:  for denoising step $t=T,\ldots,1$ do
10:     $\hat{\epsilon}\leftarrow\epsilon(\bm{z}_{t},t)$.
11:     Estimate $\hat{\bm{z}}_{t}$ using (15).
12:     $g\leftarrow\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h})$.
13:     Compute the conditional score $\bm{s}\leftarrow\zeta_{t}^{\theta}g-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}$.
14:     Sample $\varphi\sim\mathcal{N}(0,\mathbf{I})$.
15:     Compute $\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+\beta_{t}\bm{s})+\sqrt{\beta_{t}}\varphi$.
16:  end for
// Training PSD denoiser
17:  Train denoising estimators $\epsilon_{\theta}$ and $\epsilon_{\vartheta}$ for $\bm{z}$ and $\bm{h}$ following steps 3-7, respectively.
// Denoising of PSD denoiser
18:  Sample $\bm{z}_{T},\bm{h}_{T}\sim\mathcal{N}(0,\mathbf{I})$.
19:  for denoising step $t=T,\ldots,1$ do
20:     $\hat{\epsilon}_{\theta}\leftarrow\epsilon_{\theta}(\bm{z}_{t},t)$, $\hat{\epsilon}_{\vartheta}\leftarrow\epsilon_{\vartheta}(\bm{z}_{t},t)$.
21:     Estimate $\hat{\bm{z}}_{t}$ and $\hat{\bm{h}}_{t}$ using (15) and (19).
22:     $g_{\theta}\leftarrow\nabla_{\bm{z}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$, $g_{\vartheta}\leftarrow\nabla_{\bm{h}_{t}}\log p(\bm{z}^{\prime}|\bm{z}_{t},\bm{h}_{t})$.
23:     Compute the conditional score
@@ III-D Training and Implementation
24:     Sample $\varphi_{\theta},\varphi_{\vartheta}\sim\mathcal{N}(0,I)$.
25:     Compute $\bm{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}+\beta_{t}\bm{s}_{\theta})+\sqrt{\beta_{t}}\varphi_{\theta}$ and $\bm{h}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{h}_{t}+\beta_{t}\bm{s}_{\vartheta})+\sqrt{\beta_{t}}\varphi_{\vartheta}$.
26:     $\bm{h}_{t-1}\leftarrow\bm{h}_{t-1}-\alpha(||\bm{\bm{z}^{\prime}}-\bm{h}_{t-1}\bm{z}_{t-1}||_{2}+\phi||\bm{h}_{t-1}||)$.
27:  end for
// Fine-tuning process
28:  Freeze the parameters of the semantic encoder and semantic denoisers, then fed the denoised SI $\tilde{\bm{z}}$ to finetune the semantic decoder using MSE loss in (5).
Algorithm 1 Training and Implementation
Initially, we start by jointly training the semantic encoder and semantic decoder by incorporating MSE and Kullback-Leibler (KL) divergence into the loss function to better reconstruct the image. The MSE facilitates accurate reconstruction, while the KL divergence serves as a regularization term, guiding the data distribution in the latent space towards approximating a unit Gaussian distribution. The loss function is defined as the sum of MSE and KL divergence via
$\mathcal{L}=\frac{1}{N}\sum_{\bm{x}\in\bf{X}}\big{|}\big{|}\bm{x}-\tilde{\bm{x}}\big{|}\big{|}^{2}+\lambda\mathcal{L}_{\text{KL}},$ (25)
where $\lambda$ controls the significance of KL divergence relative to MSE. The KL divergence is expressed as
where $\mu_{i}$ and $\sigma_{i}$ denote the mean and standard deviation of the latent space, respectively.
Following this, we freeze the parameters of the semantic encoder and feed the latent semantic vector $\bm{z}$ into the semantic denoiser for training the SD denoiser and PSD denoiser under both known and unknown channels. Once each component is individually trained, we finalize the process by fine-tuning the entire SD-based GSC architecture in an end-to-end manner. This fine-tuning stage optimizes the collaboration between all modules, resolving any discrepancies that may arise from independently trained components. The detailed training and implementation process is presented in Algorithm 1.
@@ IV Performance Evaluation
We consider the end-to-end transmission of image tasks from a device (i.e., NVIDIA RTX 2080TI) to a server (i.e., NVIDIA A100 80GB) over a Rayleigh fading channel. The FFHQ dataset [14], consists of 70,000 high-quality human face images at a resolution of 256×256 pixels, is used as the benchmark for this evaluation. The implementation of SD-based GSC framework is built on Ubuntu 22.04 using PyTorch. For the semantic denoiser, we set the number of time steps $\mathrm{T}=1000$ and employ a linear variance scheme to determine the hyperparameters $\{\beta_{1},\beta_{2},\ldots,\beta_{\mathrm{T}}\}$. The learning rate is initialized at 0.001, and stochastic gradient descent (SGD) is used as the optimizer for the loss function.
@@ IV-A Known Channel
Fig. 2 evaluates the PSNR and the FID performance across various SNRs under known channel. As expected, all the metrics improve as the SNR increases, since better wireless channel conditions facilitate higher-quality image reconstruction. Interestingly, SD-GSC and Latent-Diff DNSC outperform ADJSCC in both metrics, due to the diffusion model’s ability in capturing wireless channel characteristics and removing noise effectively, resulting in more accurate reconstruction. Notably, SD-GSC outperforms two baselines in both metrics. Specifically, SD-GSC achieves approximately 32% and 21% PSNR improvement, and 40% and 35% FID reduction as compared to ADJSCC and Latent-Diff DNSC, respectively. This superior performance arises from the integration of the instantaneous channel gain, which enables better characterization of channel gain and mitigation of distortion caused by wirless channel noise, resulting in a more controlled and guided image generation than the uncontrolled generation process of DDPM used in Latent-Diff DNSC.
@@ IV-B Unkown Channel
To validate the effectiveness of our proposed PSD-GSC for channel estimation as well as efficient image transmission, we also compare our proposed PSD-GSC with MSD-GSC and SD-GSC. Fig. 3 presents the PSNR and FID performance for various SNRs under unknown channel. We can see both PSD-GSC and MSD-GSC experience degradation in both metrics compared to SD-GSC with known channel gain $h$. Specifically, our proposed PSD-GSC and MSD-GSC experience degradation in PSNR by approximately 1.7% and 9.0%, and improvement in FID by 4.3% and 25.5%, respectively. This is because channel estimation errors result in inaccurate channel gain inputs of the diffusion model, which decreases its denoising and high-quality image reconstruction capabilities. Importantly, PSD-GSC outperforms MSD-GSC in both metrics, improving PSNR by 8% and reducing FID by 17%, respectively. This is due to the probabilistic modeling-based diffusion model used in PSD-GSC can well estimate nonlinear channels and capture the complex properties of wireless channels, for better image reconstruction.
@@ V Conclusion
In this paper, we developed a stable diffusion-based goal-oriented semantic communication framework (SD-GSC) for efficient image transmission over Rayleigh fading channels. The main goal is to reduce the bandwidth requirement while ensuring high-quality image reconstruction at the receiver. To this end, we first designed a semantic autoencoder to extract semantic information (SI) to reduce the transmitted data size while accurately reconstructing the image from the received SI. To combat the impact of channel noise on the transmitted SI, we designed an SD-based semantic denoiser to better characterize and mitigate the channel noise, leading to significant improvements in both PSNR and FID metrics compared to the state-of-the-art baselines under known channel. For unknown channel scenarios, we further designed a parallel SD-GSC (PSD-GSC) to jointly learn the channel gain and denoise the received SI, showcasing the capability of our proposed PSD-GSC in channel estimation and efficient image transmission.